# Prompt Injection attack against LLM-integrated Applications

### v2

@@ 7.1 Defenses
While there are currently no systematic techniques established to prevent prompt injection in LLM-integrated applications, various strategies have been empirically proposed to mitigate this challenge [22, 46]. (1) Instruction Defense [46] employs a method of appending specific instructions to the prompt in order to alert the model about the subsequent content. (2) Post-Prompting [47] posits an approach where the user’s input is positioned before the prompt. (3) Random Sequence Enclosure [49] provides a security measure by enclosing the user’s input between two randomly generated character sequences. (4) Sandwich Defense [50] incorporates the user’s input within two prompts to enhance security. (5) XML Tagging [52] offers a particularly robust solution when implemented with XML+escape, by encapsulating the user’s input within XML tags, such as <user_input>. (6) Lastly, Separate LLM Evaluation [51] distinguishes potentially adversarial prompts using a distinctly prompted LLM, thus providing an additional layer of security.
@@ 8.1 LLM Security and Relevant Attacks
LLM Hallucination. Since LLMs are trained on vast crawled datasets, they have been shown to carry potential risks of generating contentious or biased content, or even perpetuating hate speech and stereotypes [8, 62, 34, 35, 17]. This phenomena is referred to as hallucination. Despite mechanisms (e.g., RLHF [54, 64]) have been introduced to enhance the robustness and reliability of the LLM outputs, there is still non-negligible risks from the target attacks.
LLM Jailbreak. Jailbreak [58, 55, 60, 33] involves eliciting model-generated content that divulges training data specifics, which can lead to serious privacy breaches, particularly when training data include sensitive or private information. Specifically, it is noticed that the content filtering can be circumvented shortly after the release of ChatGPT through jailbreak [15, 41], which typically involves hypothetical situations or simulations [58] to bypass the model restrictions. Adversaries can leverage jailbreak to abuse the model for harmful information generation.
Prompt Injection. Prompt injection [20, 44, 6] overrides an LLM’s original prompt and directs it to follow malicious instructions. This can lead to disruptive outcomes such as erroneous advice or unauthorized disclosure of sensitive information. From a broader view, backdoor [7, 68, 36] and model hijacking [56, 61] can be classified under this type of attack. Perez et al. [44] revealed GPT-3 and its dependent applications are susceptible to prompt injection attacks, which commandeer the model’s initial objective or expose the application’s original prompts. Compared to their work, we systematically explore the strategies and prompt patterns that can trigger the attack in a wider range of real-world applications.
@@ 8.2 LLM Augmentation
There is ongoing research focusing on the enhancement of LLMs to improve their operational capabilities [11, 42, 57, 59, 28, 26, 53]. An approach named Toolformer [57] demonstrates that LLMs can be trained to generate API calls, determining which APIs to use and the appropriate arguments to pass. Yao et al. [66] proposed ReAct that equips LLMs with task-specific actions and verbal reasoning based on environmental observations. There is also a shift in focus from simply integrating LLMs into applications, towards creating more autonomous systems that can independently outline solutions to tasks and interact with other APIs or models [9, 31, 30, 29, 12, 65]. An example of such a project is Auto-GPT [19], an open-source initiative capable of self-prompting to complete tasks. Another instance is Generative Agents [43] which is LLM-backed interative software to simulate human behaviors.

### v3

@@ 7.1 Defenses
While there are currently no systematic techniques established to prevent prompt injection in LLM-integrated applications, various strategies have been empirically proposed to mitigate this challenge [PromptEngineeringGuide, instructiondef]. (1) Instruction Defense [instructiondef] employs a method of appending specific instructions to the prompt in order to alert the model about the subsequent content. (2) Post-Prompting [postprompting] posits an approach where the user’s input is positioned before the prompt. (3) Random Sequence Enclosure [RSE] provides a security measure by enclosing the user’s input between two randomly generated character sequences. (4) Sandwich Defense [SandwichDefense] incorporates the user’s input within two prompts to enhance security. (5) XML Tagging [XMLTagging] offers a particularly robust solution when implemented with XML+escape, by encapsulating the user’s input within XML tags, such as <user_input>. (6) Lastly, Separate LLM Evaluation [SeparateLLME] distinguishes potentially adversarial prompts using a distinctly prompted LLM, thus providing an additional layer of security.
@@ 8.1 LLM Security and Relevant Attacks
LLM Hallucination. Since LLMs are trained on vast crawled datasets, they have been shown to carry potential risks of generating contentious or biased content, or even perpetuating hate speech and stereotypes [DBLP:conf/fat/BenderGMS21, sun2022contrastive, manakul2023selfcheckgpt, mckenna2023sources, DBLP:conf/emnlp/GehmanGSCS20]. This phenomena is referred to as hallucination. Despite mechanisms (e.g., RLHF [RLHF_news, wolf2023fundamental]) have been introduced to enhance the robustness and reliability of the LLM outputs, there is still non-negligible risks from the target attacks.
LLM Jailbreak. Jailbreak [shanahan2023role, rao2023tricking, DBLP:conf/ccs/Si0BCSZ022, liu2023jailbreaking] involves eliciting model-generated content that divulges training data specifics, which can lead to serious privacy breaches, particularly when training data include sensitive or private information. Specifically, it is noticed that the content filtering can be circumvented shortly after the release of ChatGPT through jailbreak [jailbreak_gpt1, jailbreak_gpt2], which typically involves hypothetical situations or simulations [shanahan2023role] to bypass the model restrictions. Adversaries can leverage jailbreak to abuse the model for harmful information generation.
Prompt Injection. Prompt injection [greshake2023youve, perez2022ignore, apruzzese2023realgradients] overrides an LLM’s original prompt and directs it to follow malicious instructions. This can lead to disruptive outcomes such as erroneous advice or unauthorized disclosure of sensitive information. From a broader view, backdoor [DBLP:conf/sp/BagdasaryanS22, DBLP:conf/emnlp/ZhangLM0022, mei2023notable] and model hijacking [DBLP:conf/ndss/00010022, si2023two] can be classified under this type of attack. Perez et al. [perez2022ignore] revealed GPT-3 and its dependent applications are susceptible to prompt injection attacks, which commandeer the model’s initial objective or expose the application’s original prompts. Compared to their work, we systematically explore the strategies and prompt patterns that can trigger the attack in a wider range of real-world applications.
@@ 8.2 LLM Augmentation
There is ongoing research focusing on the enhancement of LLMs to improve their operational capabilities [cai2023large, paranjape2023art, schick2023toolformer, shen2023hugginggpt, kim2023language, hao2023toolkengpt, qian2023creator]. An approach named Toolformer [schick2023toolformer] demonstrates that LLMs can be trained to generate API calls, determining which APIs to use and the appropriate arguments to pass. Yao et al. [yao2022react] proposed ReAct that equips LLMs with task-specific actions and verbal reasoning based on environmental observations. There is also a shift in focus from simply integrating LLMs into applications, towards creating more autonomous systems that can independently outline solutions to tasks and interact with other APIs or models [boiko2023emergent, liu2023chatgpt, liang2023taskmatrix, li2023api, cai2023low, xu2023tool]. An example of such a project is Auto-GPT [autogpt], an open-source initiative capable of self-prompting to complete tasks. Another instance is Generative Agents [park2023generative] which is LLM-backed interative software to simulate human behaviors.